6x+2=3y-4
Find x. How do I solve this?
Thank you.

Answer by Alan3354(69443)

6x+2=3y-4
Find x. How do I solve this?
--------------
I found it, it's on the left side next to the 6.
=========
Ok, let's get serious.
You can solve for x in terms of y.
get the x on one side of the = by itself.
----
6x+2 = 3y + 4
Subtract 2
6x = 3y + 2
Divide by 6
x = (3y/6) + (2/6)
or, x = (3y+2)/6
x = y/2 + 1/3
Either is acceptable